40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 4 04:51:14 2015, diego, Update, LSC, CARM Transition to REFL11 using CM_SLOW Path CARM_3Feb2015_CarmBwasCMslow_CarmAwasLiveALS.pdfCARM_3Feb2015_CarmBwasCMslow_CarmAwasLiveALS_Coh.pdfGlitch_in_CARM_and_PRCL_3Feb2015.pngslowCM_04-02-2015_042805.png
    Reply  Thu Feb 5 04:35:14 2015, diego, Update, LSC, CARM Transition to REFL11 using CM_SLOW Path 4Feb2015_Transition_CARM_REFL11_CM_SLOW_AP_18.png
       Reply  Fri Feb 6 03:21:17 2015, diego, Update, LSC, CARM Transition to REFL11 using CM_SLOW Path 
Message ID: 10971     Entry time: Wed Feb 4 04:51:14 2015     Reply to this: 10979
Author: diego 
Type: Update 
Category: LSC 
Subject: CARM Transition to REFL11 using CM_SLOW Path 

[Diego, Jenne, Eric]

Tonight we kept on following our current strategy for locking the PRFPMI:

  • the first few tries were pretty unsuccessful: the PRMI lock wasn't much stable, and we never managed to reduce CARM offset to zero before losing lock;
  • then we did some usual manteinance: we fixed the X arm green beatnote, fixed the phase tracker and given much attention to ASS alignment, since the X arm wasn't doing great;
  • the last few locks were consintently better: we managed to get to CARM offset zero "easily", but the arm power was not very high (~8);
  • then we tried to transition CARM to REFL11, both with the usual configuration and using CM_SLOW, using REFL11 as input for the Common Mode Board;
    • with the usual configuration, we lost lock right after the transition, because of MICH hitting the rail;
    • we did a very smooth CARM transition directly to REFL11 on the CM_SLOW path; we managed to take a spectrum with the SR785, but we couldn't take any more measurements before losing lock because of some weird glitch, as we can see from the lockloss plot;
  • another thing that helped tonight was changing the ELP of the MICH filter bank: it went from 4th order to 6th order, and from 40 dB suppression to 60 dB suppression;

both of the last two locks, the most stable ones (one transition to usual REFL11 and one transition to "CM_SLOW" REFL11) were acquired actuating on MC2;

 


EDITs by JCD:  At least one of the times that we lost PRMI lock (although kept CARM and DARM lock on ALS), was due to MICH hitting the rail, even after we increased the limiter to 15,000 counts.


 Here is the transfer function between CARM ALS (CARM_IN1) and REFL11 coming through the CM board (CARM_B), just before we transitioned over.  Coherence was taken simultaneously as usual, I just printed it to another sheet.

CARM_3Feb2015_CarmBwasCMslow_CarmAwasLiveALS.pdf

CARM_3Feb2015_CarmBwasCMslow_CarmAwasLiveALS_Coh.pdf


Here is the lockloss plot for the very last lockloss.  This is the time that we were sitting on REFL11 coming through the CM_SLOW path.  A DTT transfer function measurement was in progress (you can see the sine wave in the CARM input and output data), but I think we actually lost lock due to whatever this glitch was near the right side of the plots.  This isn't something that I've seen in our lockloss plots before.  I'm not sure if it's coming from REFL11, or the CM board, or something else.  We know that the CM board gives glitches when we are changing gain settings, but that was not happening at this time.


Q: Here's the SR785 TF of CARM locked through CM board, but still only digital control; nothing exciting. Excitation amplitude was only 1mV, which explains the noisy profile. 

Attachment 1: CARM_3Feb2015_CarmBwasCMslow_CarmAwasLiveALS.pdf  13 kB  | Hide | Hide all | Show all
CARM_3Feb2015_CarmBwasCMslow_CarmAwasLiveALS.pdf
Attachment 2: CARM_3Feb2015_CarmBwasCMslow_CarmAwasLiveALS_Coh.pdf  9 kB  | Hide | Hide all | Show all
CARM_3Feb2015_CarmBwasCMslow_CarmAwasLiveALS_Coh.pdf
Attachment 3: Glitch_in_CARM_and_PRCL_3Feb2015.png  409 kB  | Hide | Hide all | Show all
Glitch_in_CARM_and_PRCL_3Feb2015.png
Attachment 4: slowCM_04-02-2015_042805.png  149 kB  | Show | Hide all | Show all
ELOG V3.1.3-